6+3x5-4 to the power of 2.
87.93 divided 1.9 (to 2 decimal places).

Answer by checkley77(12569)

6+3x5-4 to the power of 2.
(6-4+3x^5)
(2+3x^5)^2=4+12x^5+9x^10 ans
----------------------------------------
87.93 divided 1.9 (to 2 decimal places).
87.93/1.9=46.28 ans.